How Do you find the length and width of a rectangle with a perimeter of 48 inches?

Answer by Earlsdon(6294)

With great difficulty, unless more information is provided!
Here's why:
For a rectangle, the perimeter, P is given by:
P = 2(L+W) where L = length and W = width. Substitute P = 48 inches.
48 = 2(L+W) Divide both sides by 2.
24 = L+W So, what two numbers, when added together will give 24?
1+23 = 24
2+22 = 24
3+21 = 24
4+20 = 24
5+19 = 24
6+18 = 24
7+17 = 24
8+16 = 24
9+15 = 24
10+14 = 24 etc, etc.
So, which of these is the correct answer?
You can't really say unless you have more information.